Yor are given the following information about an economy:
a. Gross private domestic investment =40
b. Goverment purchases of goods and services =30
c. Gross national product (GNP) =200
d. Current account balance =-20
e Taxes =60
f. Government transfer payment to the domesti private sector =25
g. Interest payments form the govment to the domestic private sector =15
h. (assume all intersest payments by the government goes to the domestic household)
i. Factor income received form the rest of the world =7
j. Factor payments made to rest or world =9
Find the following, assuming theat government investment is zero:
1. Consumption
2. GDP
3. Net factor payment from abroad
4. Net eport
5. Private saving
6. Government saving
7. National saving
1
Expert's answer
2019-07-08T10:20:06-0400

1. Consumption = 30+20+25+15+7 = 97

2. GDP = 40+30+60 + 25 + 15+9 = 179

3. Net factor payment from abroad = 200-179=21

4. Net export = Export - Import (no data)

5. Private saving = 40

6. Government saving = 60-30-25-15 = -10

7. National saving = 40 -10 =30